# Img-Diff: Contrastive Data Synthesis for Multimodal Large Language Models

### v2

@@ 3.2. Image Pairs Generation
We start by obtaining 118K image captions from MS COCO (Lin et al., 2014), which are descriptions biased towards real photos. Then, we use the LLM Vicuna-1.5-13B(Chiang et al., 2023) to perform object replacement in the captions. The prompt used is “Here is a sentence: ‘INPUT’. Please only replace one of the objects in this sentence with another object.” Here, INPUT refers to the original caption, and the answers from the LLM are the new captions. Finally, based on the caption pairs, we use the text-to-image generative model Stable-Diffusion-XL and the image editing technique Prompt-to-Prompt(Hertz et al., 2022) to generate image pairs with only a few objects replaced.
3.3. Difference Area Generator
3.3.1. Overview
@@
The Difference Area Generator aims to identify the locations of object differences between the image pair. Although object detection models are capable of identifying objects inside images, the range of object categories is quite limited (Zhang et al., 2022; Jocher et al., 2023). Therefore, to increase the number of detectable object categories and enhance dataset diversity, we developed the Difference Area Generator based on segmentation and image similarity comparisons. The process is illustrated in Figure 3.
@@
In this phase, we first use the Image Similarity Filter to obtain image pairs with high similarity but not completely identical and use FastSAM (Zhao et al., 2023) to perform image segmentation on each image. Next, we crop the images based on the bounding box information obtained from segmentation and use the Image-text Matching Filter to filter the cropped sub-images for the presence of valid objects. Finally, we use the Difference Detector to determine whether there are indeed differences between the bounding box regions of the image pairs and perform IoU (Intersection over Union) filtering to remove overlapping bounding boxes, ultimately obtaining valid bounding box information.
Figure 4. An overview of the Difference Captions Generator and its two stages.
[…]
An overview of the Difference Captions Generator and its two stages.
3.3.2. Image Similarity Filter
@@
The Image Similarity Filter aims to filter image pairs based on the degree of similarity. The module first uses CLIP (Radford et al., 2021) to extract image features from each image in the image pair and then calculates the cosine similarity. If their cosine similarity falls within the pre-set threshold, the image pair will be considered valid. We use the Image Similarity Filter twice in the Difference Area Generator. At the beginning stage, before using FastSAM for segmentation, we use the module to ensure that the image pairs are highly similar but not exactly the same. In the Difference Detector stage, after cropping sub-images based on the bounding box information, we use the module to filter the sub-image pairs and keep only the differing ones.
3.3.3. Image-text Matching Filter
@@
The Image-text Matching Filter determines whether an image contains valid objects (i.e., the replaced or replacing objects). This module first uses BLIP (Li et al., 2022) to extract image features, which are then compared with textual features of object names. When the image-text matching score falls within the pre-set threshold, we consider the image to contain valid objects. In the mid-stage of the Difference Area Generator, after performing sub-image cropping based on the bounding box information, we use the module to determine whether these sub-images contain valid objects and get valid bounding boxes.
3.3.4. Difference Detector
@@
The Difference Detector is used in the final stage of the Difference Area Generator. This module determines whether there are differences between the bounding box regions of the image pair. For a given bounding box, we first crop two sub-images from both images A and B based on it. These two sub-images are then filtered through the Image Similarity Filter and the bounding box is considered effective only if the difference is significant enough. After processing all bounding boxes, we use the IoU method to filter out overlapping bounding boxes. Only the bounding boxes with a higher degree of difference are retained. The remaining bounding boxes are the effective bounding boxes ultimately outputted by the Difference Area Generator.
3.4. Difference Captions Generator
3.4.1. Overview
@@
After obtaining the valid bounding box regions, we use the Difference Captions Generator to generate difference captions about the content of these regions (with each round of the process focusing on only one bounding box in one image pair). The reason our difference captions only target specific regions is that an image pair can contain multiple object differences and a single difference caption cannot fully capture all of them. Therefore, we highlight specific regions with red boxes and provide targeted difference captions to ensure greater accuracy.
@@
The module consists of two stages: the first stage generates content captions for the bounding box regions and then filters the bounding boxes with these generated captions using the Image-text Matching Filter and the Captions Similarity Filter. The second stage uses the content captions and the images highlighted with red boxes to generate difference captions. The overview process is shown in Figure 4.
3.4.2. Stage1: Object Labeling & Filter
@@
In Stage 1, for each image pair, we first select N bounding box regions with the lowest similarity between images (N is set to 5 in this project) as candidate regions. Then, for each bounding box, we use the MLLM LLaVA-NEXT (Liu et al., 2024b) to describe its corresponding regions and then apply two filtering processes: the first filter is the Image-text Matching Filter, which checks whether the content of the regions matches the captions; the second filter is the Captions Similarity Filter, which assesses whether there are differences between the two captions. Once the filtering is complete, we obtain valid bounding boxes and captions for subsequent difference captioning.
3.4.3. Captions Similarity Filter
@@
The Captions Similarity Filter determines whether the two captions corresponding to the same bounding box coordinate are different. We use CLIP to obtain text features and calculate the cosine similarity between them. When the score is low enough, we consider the two captions to be different.
3.4.4. Stage2: Difference Captions Generating
@@
In Stage 2, for each valid bounding box of each image pair, we first draw two red boxes into the images based on the bounding box information, highlighting the differences for easier localization. Then, we provide the MLLM LLaVA-NEXT with the captions of the bounding box regions and instruct it to describe the differences based on the content captions and the highlighted images. Finally, we can obtain the difference captions for the bounding box of the image pair.
3.5. Data Statistics
@@
In a nutshell, we generate 118K pairs of similar images using captions from MSCOCO and employ the Image Similarity Filter to get 38,533 highly similar but not identical image pairs. Then, we use the Difference Area Generator to filter and produce 117,779 pieces of valid bounding box information (with a maximum of 5 valid bounding boxes per image pair). Finally, we employ the Difference Captions Generator to filter and generate 12,688 high-quality “object replacement” instances.
4. Evaluation on Models trained with Img-Diff
4.1. Evaluation Settings
@@
To evaluate the effectiveness of our Img-Diff dataset, we use it to finetune the SOTA MLLMs, specifically LLaVA-1.5 (Liu et al., 2023b) and MGM (Li et al., 2023d), then evaluate them on extensive benchmarks commonly used for image difference and MLLMs. Specifically, the image difference benchmarks include the MMVP (Tong et al., 2024), the Spot-the-Diff (Jhamtani and Berg-Kirkpatrick, 2018), and the Image-Edit-Request (Tan et al., 2019).
@@
Regarding the finetuning strategy, we mix our data with the original visual instruction tuning dataset of LLaVA-1.5 and MGM respectively and conduct finetuning to get the finetuned MLLMs. For the Spot-the-Diff and Image-Edit-Request benchmarks, since they contain specific training splits, we further fine-tune MLLMs for an additional 2 epochs using only these benchmarks’ training data.
In the tables, “RP” represents “object replacement” data. For the image difference benchmarks, we select various SOTA models for comparison based on the original papers of these benchmarks and recent research on image difference captioning models.
4.2. Results on the MMVP Benchmark
@@
The MMVP benchmark is designed to systematically assess the visual capabilities of MLLMs. Its construction method is highly related to differential images: it first collects CLIP-blind pairs, which have similar CLIP features but differ in image content. Then, the differences between the images are manually described and question-answer pairs are created. Hence, the questions in MMVP is highly relevant to our dataset, as both place significant emphasis on the differences between similar images. Figure 5 presents the results on the MMVP benchmark.
Figure 5. Performance comparison on the MMVP benchmark.
[…]
Performance comparison on the MMVP benchmark.
@@
In Figure 5, we can see that finetuning LLaVA-1.5-7B and MGM-7B with our “object replacement” data noticeably improves their performance on the MMVP benchmark. After finetuning, the score of LLaVA-1.5-7B surpasses that of LLaVA-1.5-13B. Furthermore, the finetuned MGM-7B show a significant improvement in score compared to the original MGM-7B, even exceeding the scores of the SOTA models GPT-4V and Gemini. This suggests that our dataset enhances the MLLM’s ability to distinguish images with similar CLIP features but different content.
4.3. Results on the Spot-the-Diff Benchmark
@@
The dataset of Spot-the-Diff comprises a collection of street view images. These images are obtained by capturing scenes from fixed surveillance cameras at different times, resulting in pairs of street view images with minor object differences. Following previous works, our finetuned MLLMs are evaluated on BLEU(Papineni et al., 2002), METEOR(Banerjee and Lavie, 2005), CIDEr-D(Vedantam et al., 2015) and ROUGE-L(Lin, 2004). Table 1 presents the results on the Spot-the-Diff benchmark.
Table 1. Results on the Spot-the-Diff benchmark.
@@
VAM(Shi et al., 2020) 10.1 12.4 38.1 31.3
IFDC(Huang et al., 2021) 8.7 11.7 37 30.2
DUDA+Aux(Hosseinzadeh and Wang, 2021) 8.1 12.5 34.5 29.9
VACC(Kim et al., 2021) 9.7 12.6 41.5 32.1
@@
Based on Table 1, it is evident that after finetuning with our “object replacement” data, both LLaVA-1.5-7B and MGM-7B achieve significant performance improvements on the Spot-the-Diff benchmark, which indicates that our dataset enhances the ability of MLLMs to detect subtle differences between similar images.
4.4. Results on Image-Editing-Request
The Image-Editing-Request benchmark focuses on image editing. Each instance in its dataset consists of an image pair (i.e., a source image and an edited image) and an editing instruction that describes the transformation from the source image to the edited image. During the evaluation, our models are required to generate transformation descriptions for these image pairs, and we then calculate the BLEU, METEOR, CIDEr-D, and ROUGE-L scores with the models’ responses and the reference answers. Table 2 presents the results of the Image-Editing-Request benchmark.
Table 2. Results on the Image-Edit-Request benchmark.
@@ 4.4. Results on Image-Editing-Request
VARD(Tu et al., 2023a) 10 14.8 35.7 39
CLIP4IDC(Guo et al., 2022) 8.2 14.6 32.2 40.4
NCT(Tu et al., 2023b) 8.1 15 34.2 38.8
BiDiff(Sun et al., 2022) 6.9 14.6 27.7 38.5
VIXEN(Black et al., 2024) 8.6 15.4 38.1 42.5
@@
Table 3. Performance comparison on 8 MLLM benchmarks.
Model VQAv2 GQA POPE MMB MMBCN $M$-Vet SQAI SEED $\triangle$
LLaVA-1.5-7B 78.5 62 85.9 64.3 58.3 30.5 66.8 58.6 -

### v3

@@ 3.2 Image Pairs Generation
We start by obtaining image captions from caption databases, which contain descriptions biased towards real photos. Then, we use an LLM to perform object replacement in the captions. The prompt used is “Here is a sentence: ‘INPUT’. Please only replace one of the objects in this sentence with another object.” Here, INPUT refers to the original caption. Next, based on the caption pairs, we use the Stable-Diffusion-XL and the Prompt-to-Prompt to generate image pairs with only few objects replaced.
Figure 4: An overview of the Difference Captions Generator and its two stages.
3.3 Difference Area Generator
3.3.1 Overview
@@
The Difference Area Generator aims to identify the locations of object differences between the two images in pairs. Although object detection models are capable of detecting objects in images, the range of object categories is quite limited [74, 28]. Therefore, to increase the number of detectable object categories and enhance dataset diversity, we develop the Difference Area Generator based on segmentation and image similarity comparisons. The process is illustrated in Figure 3.
@@
We first use an Image Similarity Filter to obtain image pairs with high similarity but not identical. Next, we use the FastSAM [78] to perform image segmentation on each image. After that, we crop the images based on the bounding box information obtained from segmentation and use an Image-text Matching Filter to filter the cropped sub-images for the presence of valid objects. Finally, we use a Difference Detector to determine whether there are indeed differences between the bounding box regions of the two images and perform IoU (Intersection over Union) filtering to remove overlapping bounding boxes, ultimately obtaining valid bounding box information.
3.3.2 Image Similarity Filter
@@
The Image Similarity Filter aims to filter image pairs based on the degree of similarity. The module first uses CLIP [55] to extract image features from each image in pairs and then calculates the cosine similarity score. If their score falls within the pre-set threshold, the image pair will be considered valid. Specifically, we use the module twice in the Difference Area Generator. Before the segmentation, we use the module to ensure that the images in pairs are highly similar but not the same. In the Difference Detector stage, after cropping, we use the module to filter the sub-image pairs and keep only the differing ones.
3.3.3 Image-text Matching Filter
@@
The Image-text Matching Filter determines whether an image contains valid objects (i.e. the replaced or replacing objects). This module first uses BLIP [35] to extract image features, which are then compared with textual features of object names. When the image-text matching score falls within the pre-set threshold, we consider the image to contain valid objects. In the mid-stage of the Difference Area Generator, after cropping, we use the module to determine whether these sub-images contain valid objects and get valid bounding boxes.
3.3.4 Difference Detector
@@
The Difference Detector determines whether there are differences between the bounding box regions of the two images. Based on a given bounding box, we first crop two sub-images from both image A and image B. The two sub-images are then filtered through the Image Similarity Filter and the bounding box is considered effective only if the difference is significant enough. After processing all bounding boxes, we use the IoU method to filter out overlapping bounding boxes. Only the bounding boxes with a higher degree of difference are retained.
3.4 Difference Captions Generator
3.4.1 Overview
@@
After obtaining the valid bounding box regions, we use the Difference Captions Generator to generate descriptions for the differences inside these areas (with each round of the process focusing on only one bounding box in one image pair). Evidently, an image pair can contain multiple object differences but a single difference caption cannot fully capture all of them. Therefore, we highlight specific regions with red boxes and provide targeted difference captions to ensure greater accuracy.
@@
The module consists of two stages: the first stage generates captions for the content in the bounding box regions and then filters the bounding boxes using an Image-text Matching Filter and a Captions Similarity Filter. The second stage uses the content captions and the images highlighted with red boxes to generate difference captions. The overview is shown in Figure 4.
3.4.2 Stage1: Object Labeling & Filtering
@@
In Stage 1, for each image pair, we first select N bounding box regions with the lowest similarity between images (N is set to 5 in this project) as candidate regions. Then, for each bounding box, we use the MLLM LLaVA-NEXT [44] to describe its corresponding regions and then apply two filtering processes: the first filter is an Image-text Matching Filter, which checks whether the content of the regions matches the captions; the second filter is an Captions Similarity Filter, which assesses whether there are differences between the two captions. Once the filtering is complete, we obtain valid bounding boxes and captions for subsequent difference captioning.
3.4.3 Captions Similarity Filter
@@
The Captions Similarity Filter determines whether the two captions of the same bounding box coordinates are different. We use CLIP to obtain text features and calculate the cosine similarity between them. When the score is low enough, we consider the two captions to be different.
3.4.4 Stage2: Difference Captions Generating
@@
In Stage 2, for each valid bounding box of each image pair, we first draw two red boxes into the images based on the bounding box coordinates, highlighting the differences for easier localization. Then, we provide the MLLM LLaVA-NEXT with the captions of the bounding box content and instruct it to describe the difference based on these captions and the highlighted images. Finally, we can obtain the difference caption for the highlighted area.
3.5 Data Statistics
@@
Using captions from MSCOCO [40], we generate 118K pairs of similar images. We then employ the Image Similarity Filter to get 38,533 highly similar but not identical image pairs. Next, we use the Difference Area Generator to filter and produce 117,779 pieces of valid bounding box information (with a maximum of 5 valid bounding boxes per image pair). Finally, we employ the Difference Captions Generator to filter and generate 12,688 high-quality “object replacement” instances. Our evaluations on the main page are based on this dataset.
In addition, we also use captions from the LLaVA pre-training dataset to generate 34,538 “object replacement” samples. In Section 10 of the appendix, we evaluate this dataset and discuss the relationship between data quantity, data quality, and model performance improvement, emphasizing the importance of prioritizing quality over quantity.
4 Evaluation and Main Results
4.1 Evaluation Settings
@@
To evaluate the effectiveness of the Img-Diff dataset, we use it to fine-tune SOTA MLLMs, including LLaVA-1.5-7B, MGM-7B, and InternVL2-8B on the main page, as well as InternVL2-1B and LLaVA-1.5-13B in the supplementary materials. We evaluate these models on extensive benchmarks commonly used for image difference and MLLMs. Specifically, the image difference benchmarks include MMVP [63], Spot-the-Diff [26], and Image-Edit-Request [62]. Besides, the details and results of the MLLM benchmarks are shown in Section 4.5.
@@
During our fine-tuning process, we first mix our data with MLLMs’ original visual instruction tuning data respectively. For LLaVA-1.5 and MGM, we conduct the fine-tuning anew. For InternVL2, we follow the guidelines from its official repository and perform a second fine-tuning. (To ensure a fair comparison, we perform a second fine-tuning on InternVL2 using its original fine-tuning dataset, termed as InternVL2-FT, and use it as a baseline.) Regarding the Spot-the-Diff and Image-Edit-Request benchmarks, as they contain data of training splits, we further fine-tune the fine-tuned MLLMs for an additional 2 epochs using only these benchmarks’ training data.
In the tables, “RP” represents “object replacement” data.
4.2 Results on the MMVP Benchmark
@@
The MMVP benchmark is designed to systematically assess the visual capabilities of MLLMs. Its data processing method is highly related to image difference: it first collects CLIP-blind pairs, which have similar CLIP features but differ in image content. Then, the differences between the images are manually described and question-answer pairs are created. Hence, the questions in MMVP is highly relevant to our dataset, as both place significant emphasis on the differences between similar images.
Figure 5: Performance comparison on the MMVP benchmark.
@@
As shown in Figure 5, fine-tuning MLLMs with our “object replacement” data significantly improves their performance on the MMVP benchmark. After fine-tuning, the score of LLaVA-1.5-7B exceeds that of LLaVA-1.5-13B. Furthermore, the fine-tuned MGM-7B shows a significant improvement in score compared to the original MGM-7B, even surpassing the scores of the SOTA models GPT-4V and Gemini by up to 12. Additionally, the performance of InternVL2-8B is also enhanced. These results suggest that our dataset enhances MLLMs’ abilities to distinguish images with similar CLIP features but different content.
4.3 Results on the Spot-the-Diff Benchmark
@@
The dataset of Spot-the-Diff comprises pairs of street view images that display subtle object differences. These images are obtained by capturing scenes from fixed surveillance cameras at different time.
Table 1: Performance comparison on Spot-the-Diff.
@@
VAM[59] 10.1 12.4 38.1 31.3
IFDC[23] 8.7 11.7 37 30.2
DUDA+Aux[22] 8.1 12.5 34.5 29.9
VACC[30] 9.7 12.6 41.5 32.1
@@
InternVL2-8B-FT 6.6 11.7 26.5 27.3
InternVL2-8B + RP 8.4 12.8 32.2 28.5
Following previous works, our fine-tuned MLLMs are evaluated on BLEU [51], METEOR [4], CIDEr-D [68] and ROUGE-L [39]. As shown in Table 1, after fine-tuning with our “object replacement” data, both LLaVA-1.5-7B, MGM-7B, and InternVL2-8B achieve significant performance gains on Spot-the-Diff. Compared to score increases seen in prior models, our dataset yields even greater enhancements than those from previous iterations of image difference models. These positive results indicate that our dataset can enhance the ability of MLLMs to detect subtle differences between similar images.
4.4 Results on Image-Editing-Request
The Image-Editing-Request benchmark focuses on image editing. Each instance in its dataset consists of an image pair (i.e. a source image and an edited image) and an editing instruction that describes the transformation. During the evaluation, our models are required to generate transformation descriptions for these image pairs, and we then calculate the BLEU, METEOR, CIDEr-D, and ROUGE-L scores using the models’ responses and the reference answers.
Table 2: Performance comparison on Image-Edit-Request.
@@ 4.4 Results on Image-Editing-Request
VARD[65] 10 14.8 35.7 39
CLIP4IDC[18] 8.2 14.6 32.2 40.4
NCT[66] 8.1 15 34.2 38.8
BiDiff[61] 6.9 14.6 27.7 38.5
VIXEN[5] 8.6 15.4 38.1 42.5
@@
InternVL2-8B-FT 12.4 14.1 51.5 38.9
InternVL2-8B + RP 12.5 14.2 56.0 39.4
As shown in Table 2, LLaVA-1.5-7B, MGM-7B, and InternVL2-8B originally show SOTA performance on the Image-Edit-Request benchmark. Upon incorporating our “object replacement” data for better fine-tuning, the performance of all three models improves even more, achieving new SOTA scores. This suggests that our dataset enhances MLLMs’ abilities to recognize similarities and dissimilarities in image pairs, as well as enables them to describe differences more accurately.
4.5 Results on MLLM Benchmarks
Aside from the evaluations related to image difference discrimination, we also assess the performance of our “object replacement” data in enhancing the comprehensive abilities of MLLMs. We test the fine-tuned MLLMs using commonly used MLLM benchmarks, including VQAv2 [17] and GQA [24] for assessing the comprehensive VQA capabilities; MMBench [46], MM-Vet [72], ScienceQA [47], and SEED-Bench [34] for testing perceptual and reasoning abilities; and POPE [37] for evaluating fine-grained object localization abilities. Table 3 presents the results on these MLLM benchmarks, with the $\triangle$ metric indicating the percentage improvement averaged across them.
Table 3: Performance comparison on 8 MLLM benchmarks.
Model VQAv2 GQA POPE MMB MMBCN
LLaVA-1.5-7B 78.5 62 85.9 64.3 58.3
LLaVA-1.5-7B + RP 79.3 $\uparrow$ 62.8 $\uparrow$ 86.4 $\uparrow$ 66.1 $\uparrow$ 59.8 $\uparrow$
MGM-7B 80.4 62.6 86 69.3 58.9
MGM-7B + RP 80.7 $\uparrow$ 62.7 $\uparrow$ 86.2 $\uparrow$ 68.7 $\downarrow$ 59.6 $\uparrow$
InternVL2-8B-FT 81.8 62.6 87.7 82.5 81.5
InternVL2-8B + RP 81.8 62.6 88.0 $\uparrow$ 82.7 $\uparrow$ 81.4 $\downarrow$
Model $M$-Vet SQAI SEED $\triangle$
LLaVA-1.5-7B 30.5 66.8 58.6 -